19:45 Stop and restart firefox on lbtmu03, since it was taking up to 62% memory and was slow to respond. In new firefox, start LBC and run up.

19:50 Start series of shutter tests, taking 1-5 sec exposures at each of the filters. Telescope is at horizon and there is some light in dome. Series begins with lbcb.20070916.024954.fits

next morning We continued shutter tests in the closed doem, telescope fixed, for a while and it worked - 176 images. Fog came in --- still here in the morning.]

-Olga Kuhn

-- JohnHill - 17 Sep 2007
Topic revision: r2 - 17 Sep 2007, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback